You've collected the following information from a financial web site.

52-Week Price Stock (Div) Div
Yld %
PE
Ratio
Close
Price
Net
Chg
Hi Lo
77.40        10.43          Palm Coal 0.36 2.6       6        13.90     –0.24    
56.56        34.17          Lake Lead Grp 2.29 5.6       10        41.18     –0.01    
130.93        69.50          SIR 2.00 2.2       10        88.97     3.07    

What was the highest dividend yield over the past year?

What was the lowest dividend yield over the past year?
Answer #1

Solution:

Dividend yield: dividend yield is the dividend paid divided by the current share price.

Part A) Highest Dividend Yield

It can be seen from the table above that the highest dividend yield was for   Lake Lead Grp at 5.6%.

it can be calculated by Yield = Dividend / Current share price = 2.29/41.18 = 5.6%

Part B) Lowest Dividend Yield

It can be seen from the table above that the lowest dividend yield was for SIR at 2.2%.

Yield = Dividend / Current share price = 2.2/88.97 = 2.2%